OpenStudy (anonymous):

y=3/2x+1, -5/2x-2, y=1/3x+5 Findin the slope and y-int

OpenStudy (ash2326):

y=mx +c m=slope c=intercept on y axis 1st m=3/2, c=1 2nd m=-5/2 c=-2 3rd m=1/3 c=5